OpenStudy (anonymous):

You deposit $6,700 in a money-market account that pays an annual interest rate of 3.3%. The interest is compounded quarterly. How much money will you have after 4 years?

OpenStudy (wolf1728):

total = princ * (1 + rate)^n where 'n' is number of years * # of compounding periods total = 6,700 * (1+0.00825)^16 total = 6,700 * 1.1404905494 total = 7,641.29

OpenStudy (anonymous):

thank you!

OpenStudy (wolf1728):

u r welcome